OpenStudy (anonymous):

given F(x)=3x+5 and G(x)=x^2 a. (f o g)(x) b.(g o f) (x) c.(f o g) (-x)

OpenStudy (anonymous):

First step is to make sure of their domain and range. D(F)=R, R(F)=R D(G)=R, R(G)=R/R- So, all three are possible. as R(F) is subset to D(G) and vice versa. I'll do the first one as example. \[ (f o g)(x)=f(g(x))=f(x^2)=3x^2 + 5\] i.e just substitute g(x) into x. The others are similar.

OpenStudy (anonymous):

@Shadowys so how do you do (c)?

OpenStudy (anonymous):

I'll substitute -x for x but since g(x) is an even function, g(-x)=g(x) and (f o g)(x)=(f o g) (-x)

OpenStudy (anonymous):

@Shadowys basically a and c are the same?

OpenStudy (anonymous):

Yes. note this is a special case of g(x).

OpenStudy (anonymous):

@Shadowys Thankkkss

OpenStudy (anonymous):

:)